If a recipe requires 100 grams of grated motzarella for a 25cm pepperoni pizza, how much cheese will be neede for a 40inch pizza? (the measurement given here is the diameter of the pizza.)
thank you so much for giving the time to answer ,
please answer asap. i really need it now, great thanks.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

If a recipe requires 100 grams of grated motzarella for a 25cm pepperoni pizza, how much cheese will be neede for a 40inch pizza? (the measurement given here is the diameter of the pizza.)
-----------
1 inch = 2.54 cm
40 inches = 101.6 cm
----------------
Area is a function of the square of the diameter, so the area of the 40 inch pizza is (101.6/25)^2 times the area of the 25 cm pizza
Amount = 100*(101.6/25)^2 = 1651.6096 grams
=~ 1.65 kg

Answer by stanbon(75887)

If a recipe requires 100 grams of grated motzarella for a 25cm pepperoni pizza, how much cheese will be neede for a 40inch pizza? (the measurement given here is the diameter of the pizza.)
------------------------
Area of a 25cm pizza = 25*pi sq. cm.
----
Area of a 40 INCH piaza is 40*pi sq inches.
-------------
Note: I have a feeling you did not post
this problem correctly.
Cheers,
Stan H.
========